OpenStudy (anonymous):

Add and simplify the following problem..Thanks 5/49z^3+1/35z^2

OpenStudy (saifoo.khan):

\[\frac{7z+25}{245z^3}\]

hero (hero):

Saifoo solved it earlier but forgot to provide the steps. He was trying to be a hero.

OpenStudy (saifoo.khan):

Lame joke^

OpenStudy (saifoo.khan):

@MsLorraine, If you want steps, i can provide you with those as well.

hero (hero):

It's not a joke

OpenStudy (anonymous):

@saifoo.khan....no I believe they just want the simplified answer

OpenStudy (saifoo.khan):

Thankyou! :)

OpenStudy (anonymous):

im sorry yes......show me the steps and then answer

OpenStudy (anonymous):

im preparing for a final in 4wks...and I have a long way to go....Algebra has never been my favorite subjects, but its needed in my field.

hero (hero):

I'm convinced that something is wrong with your numerator

hero (hero):

@saifoo

OpenStudy (anonymous):

Im sorry saifoo...I did not see ur answer earlier...i was using a public computer, and it kept giving me problems....is the answer 7z+25/245z^3 right?

OpenStudy (saifoo.khan):

@hero, if it's wrong prove it please. @MsLorraine, Yes Exactly!

OpenStudy (anonymous):

thank you khan and hero...u all r my saviors...because I do not want to fail math...I am going for my Masters in Healthcare Administration and I must know this math...

OpenStudy (saifoo.khan):

Best of luck with that! Do u have more problems? Feel free to post them here.

OpenStudy (anonymous):

here is another one; -63+7x/5divided by 4x-36/25=

OpenStudy (anonymous):

just a quick question saifoo.khan, are u a math major....or do u just like math?

OpenStudy (saifoo.khan):

Im a college math tutor! :)

OpenStudy (saifoo.khan):

\[\frac{35}{4}\] Answer.

OpenStudy (anonymous):

oh ok...have u considered becoming a tutor for university of phoenix...because they definitely need it....because the ones that they have, has caused me to fail...so I dont use them at all...

OpenStudy (anonymous):

But if you can guarantee ur work and u r 100%, you will be a lot of help for a lot of people at the college.

OpenStudy (saifoo.khan):

Not really, i havnt applied for teaching there, but i have been teaching UoP guys from 10 months now.

OpenStudy (saifoo.khan):

Thankyou!! :)

OpenStudy (anonymous):

Also if I receive a good grade, I will recommend you to everyone in my math class, because we need someone like you...

OpenStudy (anonymous):

Is your work guaranteed?

OpenStudy (saifoo.khan):

Awesome, you may give them my email address. here's my email, einstein.saifoo@gmail.com Email me, we can talk! :)

OpenStudy (saifoo.khan):

einstein.saifoo@gmail.com

OpenStudy (anonymous):

got it...

OpenStudy (saifoo.khan):

Cool.

OpenStudy (anonymous):

35/4 is the answer.....the problem says simplify...would that be 8.75 divided right?

OpenStudy (saifoo.khan):

Nope, that must be saying "write a fraction"

OpenStudy (anonymous):

okay, so when they ask to simplify, it doesn't always mean divide

OpenStudy (saifoo.khan):

Yes! That's the simplest form.

OpenStudy (anonymous):

okay, I will notate that...

OpenStudy (anonymous):

here is another one; x^2+8x/x^2+3x-40 divided by x/x+5=

OpenStudy (saifoo.khan):

\[\Large \frac{x^2}{x^2-25}\]

OpenStudy (saifoo.khan):

Alright.off to email. :)

OpenStudy (saifoo.khan):

Hey? I emailed you something..

OpenStudy (anonymous):

ok